Part 3: Geometry The home you ended up buying with the $125,000 mortgage is pictured hene.

media%2Fc7b%2Fc7b89905-7b20-46ef-9444-18

media%2F5e6%2F5e613352-04b6-4e12-8d2f-9d
Answer #1

Question1 of part3 exercise.

As we know that ,

Area of rectangle = length x breadth

Therefore,

Area of front door is given by

= 80×36 in2

= 2880 in2

= 2880/144 ft2(\large \because there are 144 square inches in 1 square feet i.e. 1ft2 = 144 in2 because there are 12 inches in 1 feet )

= 20 square feet

Area of window 1 is given by

= 24×30 in2

= 720 in2

= 720/144 ft2

= 5 square feet

Area of window 2 and 4 is given by

= 30x36 in2

= 1080 in2

= 1080/144 ft2

= 7.5 square feet

Area of window 3 is given by

= 12×20 in2

= 240 in2

= 240/144 ft2

=1.67 square feet

Area of window 5 and 6 is given by

= 36x20 in2

= 720 in2

= 720/144 ft2

= 5 square feet